OpenStudy (anonymous):

helppppppppppppp! Can you please help me with this? Make p the product of the formula.q=5/3+p @Mertsj

OpenStudy (mertsj):

\[q=\frac{5}{3+p}\]

OpenStudy (mertsj):

Is that the question?

OpenStudy (anonymous):

YES

OpenStudy (mertsj):

Multiply both sides by 3+p

OpenStudy (anonymous):

I DONT KNOW ANYTHING ABOUT IT

OpenStudy (mertsj):

Then why do you have this problem if you don't know anything about it?

OpenStudy (anonymous):

I WAS GIVEN TO SEE HOW HOW I KNOW

OpenStudy (anonymous):

BUT IF I HAVE AN UNDERSTANDING IT WILL BE EVEN BETTER WHEN I GET IN CLASS

OpenStudy (mertsj):

\[(3+p)\times q=\frac{5}{3+p} \times (3+p)\]

OpenStudy (mertsj):

\[3q+pq=5\]

OpenStudy (mertsj):

\[pq=5-3q\]

OpenStudy (mertsj):

\[pq \times \frac{1}{q}=(5-3q) \times\frac{1}{q}\]

OpenStudy (mertsj):

\[p=\frac{5-3q}{q}\]

OpenStudy (anonymous):

THANKS VERY VERY MUCH THE WORKINGS HELP A LOT

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

I APPRICATE IT HOW DO I GIVE U A MEDAL MY COMPUTER WILL GO OFF IN A WHILE